# Long-Form Text-to-Music Generation with Adaptive Prompts: A Case of Study in Tabletop Role-Playing Games Soundtracks

### v1

@@ 1 Introduction
Figure 1: At every 30 seconds of gameplay, Babel Bardo transcribes the players’ speeches into a text $s_{i}$ using a Speech Recognition system and uses a Large Langue Model (LLM) to map $s_{i}$ into a music description $d_{i}$ that matches the scene described by the players. This music description is given to a Text-to-Music system that generates a 30-second piece $a_{i}$ directly in the audio domain.
@@ B E D DC
Table 1: FAD for each Babel Bardo version in COTW and OSNI in contrast with Human music, i.e., the original soundtracks used by the player in both these TRPGs.
@@
In our experimental setup, we employed the YouTube API as our Speech Recognition system, extracting the transcriptions from the TRPG campaigns. We used Ollama 3.1, with 70B parameters, as Babel Bardo’s LLM for generating the music descriptions. As our text-to-music model, we used MusicGen [2] large, with 1.3B parameters. For every transcription $s_{i}$, we generate a 30-second long audio signal $a_{i}$, which is the maximum length MusicGen supports. We used the VGGish model for computing FAD scores. We computed the KL-Divergence with the PaSST [22] classifier, which was pre-trained on the AudioSet dataset [23]. We fine-tuned PaSST on the MTG-Jamendo dataset [24] to have a model more semantically suited for soundtrack classification (e.g., mood, genre, instrumentation). We fine-tuned PaSST for a single epoch with a learning rate of $10^{-4}$.
@@
Table 3: Mean/Standard Deviation of transition KLD for each Babel Bardo version in both both COTW and OSNI.

### v2

@@ 1 Introduction
Figure 1: At every 30 seconds of gameplay, Babel Bardo transcribes the players’ speeches into a text $s_{i}$ using a Speech Recognition system and uses a Large Language Model (LLM) to map $s_{i}$ into a music description $d_{i}$ that matches the scene described by the players. This music description is given to a Text-to-Music system that generates a 30-second piece $a_{i}$ directly in the audio domain.
@@ B E D DC
Table 1: FAD for each Babel Bardo version in COTW and OSNI in contrast with Human music, i.e., the original soundtracks used by the players in both these TRPGs.
@@
In our experimental setup, we employed the YouTube API as our Speech Recognition system, extracting the transcriptions from the TRPG campaigns. We used Ollama 3.1, with 70B parameters, as Babel Bardo’s LLM for generating the music descriptions. As our text-to-music model, we used MusicGen [2] large, with 3.3B parameters. For every transcription $s_{i}$, we generate a 30-second long audio signal $a_{i}$, which is the maximum length MusicGen supports. We used the VGGish model for computing FAD scores. We computed the KL-Divergence with the PaSST [22] classifier, which was pre-trained on the AudioSet dataset [23]. We fine-tuned PaSST on the MTG-Jamendo dataset [24] to have a model more semantically suited for soundtrack classification (e.g., mood, genre, instrumentation). We fine-tuned PaSST for a single epoch with a learning rate of $10^{-4}$.
@@
Table 3: Mean/Standard Deviation of transition KLD for each Babel Bardo version in both COTW and OSNI.